OpenStudy (hanner_b_nanner):

I need ALOT of help..Fan&&Medal I'm doing Quadratic Equations and I'm totally lost. Solve x^2-3x=-8 For some reason I got 1+/- i sqrt(95) over 2

OpenStudy (hanner_b_nanner):

@mathmath333

OpenStudy (mathmath333):

its incorrrect i rhink

OpenStudy (mathmath333):

*think lol

OpenStudy (hanner_b_nanner):

It is.. all my answers are either 3 or -3 +/- i sqrt(29) or (23) over 2... :(

OpenStudy (mathmath333):

\(\large\tt \color{black}{ax^2+bx+c=0}\) \(\Huge\tt \color{black}{x=\dfrac{-b\pm\sqrt {b^2-4ac}}{2}}\)

OpenStudy (hanner_b_nanner):

 I know that's not right... grrrrr lol